At the fair, the chickens and the rabbits were placed under the same tent. The chickens and the rabbits have 360 legs. If there were 105 animals how many were rabbits ?
Answer by Edwin McCravy(20060)

At the fair, the chickens and the rabbits were placed under the same tent. The chickens and the rabbits have 360 legs. If there were 105 animals how many were rabbits ?
If they were all chickens, there would be only 2×105 or 210 legs.
So the extra 360-210 or 150 legs belong to rabbits.  That's 150÷2
or 75 pairs of hind legs, so there are 75 rabbits.

You can also do it with algebra:

c + r = 105
2c + 4r = 360 

c = 105 - r

2(105-r) + 4r = 360
210 - 2r + 4r = 360
           2r = 150
            r = 75

Edwin